OpenStudy (anonymous):

Complete the square to write c(x) = x2 - 16x + 84 in vertex form.

OpenStudy (anonymous):

First step: Rewrite your equation like this: c(x) = (x2 - 16x + ____) + 84 Now we need to figure out what we want to be in that ____

OpenStudy (anonymous):

b/2 quantity squared

OpenStudy (anonymous):

ok.. lets wait for a response here.

OpenStudy (anonymous):

wouldn't you put 16 in the blank?

OpenStudy (anonymous):

and then add 16 to the other side?

OpenStudy (anonymous):

no. like ehuman said you want (b/2)^2 in there. b = 16 in our example. so take 16 / 2 and square the result. what do you get?

OpenStudy (anonymous):

because 16/2 = 8 8 x 8 = 64!

OpenStudy (anonymous):

yes. put a 64 in the ___. now, how do we compensate for adding that 64?

OpenStudy (anonymous):

you add it to the opposite side? or do you add it to the 84?

OpenStudy (anonymous):

subtract it. we are working on the same side of the equation...so to compensate for adding something, on the same side just subtract the same number. c(x) = (x2 - 16x + 64) + 84 - 64

OpenStudy (anonymous):

okay, so let me try and do the rest of the work and I will post it:)

OpenStudy (anonymous):

alright. good luck

OpenStudy (anonymous):

\[c(x) = (x^2 - 16x + 64) + 84 - 20\] \[c(x) = (x^2 - 16x + 64) + 20\] \[c(x) = (x - 8)(x - 8) + 20\] \[c(x) = (x - 8)^2 + 20\] ?? @BangkokGarrett

OpenStudy (anonymous):

the end there should be + 64 + 84 - 20 is + 64